OpenStudy (anonymous):

George and Carmren went on a bicycle trip. They took a bus to their starting point, and then biked the rest. They have traveled 300 kilometers in total, and they biked 70 kilometers more than they were bussed. Let x= kilometer traveled by bike and y=kilometer travelled by bus Find out how many kilometers they traveled by bike.

OpenStudy (anonymous):

So, x + y = 300 x - y = 70

OpenStudy (anonymous):

Solve them using substitution or something.

OpenStudy (anonymous):

x=(300/2)+(70/2) x=150+35 what is 150+35 @mathsims

OpenStudy (anonymous):

@mathsims are u there